# Reinforcement Learning with Dynamic Multi-Reward Weighting for Multi-Style Controllable Generation

### v1

@@ 4.3 Configurations of (Custom) PPLM
Reward Formulation Negative normal-$\uparrow$ Informal normal-$\uparrow$ Neg & Inf normal-$\uparrow$ PPL normal-$\downarrow$ Bigram Dup normal-$\downarrow$
@@
 GradWt Softmax 65.90 76.70 60.25 31.46 0.1665
Table 2: Comparison of reward formulations for the Negative + Informal style combinations. We find that GradWt Softmax rewards show the best control over the style combination while maintaining generation quality.
Src Dataset Prompt
SemEval 2017
Its nice to stay
TweetEval
@user aha but white
Rotten Tomatoes
One of the most
Wikipedia
Gorman Park (or Amelia
Table 3: Example prompts used in our pipeline.
@@ 4.4 Training and Evaluation Prompt Data
We utilize a collection of prompts for use during training and inference. Our prompts during training are based on items from the training set of the following datasets: SemEval 2017 emotions (Rosenthal et al., 2017), TweetEval (Barbieri et al., 2020), Rotten Tomatoes (Pang and Lee, 2005), and Wikipedia datasest (Wikimedia, ). We choose these datasets for their variety of domains and styles. For evaluation, we prompt models with 500 randomly selected items from the test sets of these datasets, prompting with the first four words of the item. Examples are in Table 3.
4.5 Evaluation Metrics
We evaluate the generations from each model based on two general principles: first, how often the generations adhere to the target styles, as defined by the style discriminators used during training. To measure adherence to the target styles, we count the proportion of the generations that are classified by the discriminators as having the target style. We also count the proportion of the generations that have both target styles, as a model may output relatively high percentages of a target style but unsuccessfully combine them in the same generation.
Second, we evaluate whether the generations maintain the linguistic quality of the original model. Consider a model fine-tuned to produce positive generations: it could simply respond to every prompt with “This is great, I love it!” and force every generation to be classified as positive, but the overall language quality would suffer. To evaluate whether the generations maintain linguistic quality, we measure the average perplexity of the generations as well as their repetitiveness (specifically, how often duplicate bigrams appear across all generations from a particular model).
@@ 5 Results
In this section we describe our experiments for assessing the different multi-style reward formulations and the overall performance of our models for various multi-style control settings. The questions we investigate are:
  • How can we most effectively combine signals from separate style discriminators into a reward function? (Section 5.1)
  • How often do two- and three-style models express the target styles in their generation, and how fluent are these generations? (Section 5.25.3)
  • Which style combinations are most difficult to learn? (Section 5.4)
For more perspective on these results, we also conduct experiments using a decoding-time PPLM approach and zero-shot prompting with GPT-4.
Model Base (Llama2-7B) FineTune with GradWt (ours) PPLM
@@
Table 4: Automatic target style accuracy evaluation of 1- and 2-style models for RL and PPLM. (“Both” indicates the portion of generations containing both the target sentiment and formality.) All models (Base, FineTune with GradWt, PPLM) use Llama2-7B as the base model for fair comparison. Compare style frequencies to those in the base LLaMA2 generations to see the effect of the other two approaches.
Reward Hacking (Negative + Informal)
You have no affinity for this community. I don’t know what to say, I don’t know what to do.
Violent, vulgar, obscenely disrespectful. I’m just going to list the stuff, you know, for the people who are…I don’t know, I don’t work on
Ugh. Our Pho looks so bad. I don’t know how to say this, but, it does
Table 5: Models can obtain high rewards by repeating the same phrases (in this case, “I don’t know”). Our hyperparameter searches prioritize models without such reward hacking, even if they obtain lower rewards.
5.1 Multi-Reward Control Formulation Evaluation
We train several versions of models with Informal-Negative and Formal-Negative target styles to investigate the efficacy of the reward formulations outlined in Sec 3.2. For a summary of our study on Informal-Negative see Table 2; full results are in Appendix Table 8. We find that for softmax rewards, the models make only small improvements on controlling styles. We hypothesize that a contributing factor may be poor confidence calibration of our discriminators’ softmax scores. To correct this, we implement a calibrated softmax reward using the confidence calibration technique proposed in Guo et al. (2017). This results in the Expected Calibration Error (ECE) for sentiment and formality discriminators decreases from $0.133$ to $0.016$ and $0.267$ to $0.111$ respectively. These calibrated softmax rewards do offer some improvements relative to the un-calibrated softmax rewards, but improvements are modest.
Reward formulations using raw logits, on the other hand, tend to control the styles better at the cost of fluency, as measured by the average perplexity of the generations. We also anecdotally observed that the logit-reward models often developed highly repetitive reward hacking strategies when the KL coeffecient was not sufficiently high; see examples in Table 5.
Both the binary reward approach and dynamic gradient weighting (GradWt) approach produces sizable improvements over the softmax or logit based rewards while maintaining generation quality. The GradWt approach has a slight edge overall: it outperforms the binary approach on all counts for the Negative-Informal combination ($60.25\%$ of GradWt generations contain both styles versus $56.80\%$ for binary). Generations from the binary approach are also more repetitive than those from GradWt. Negative-Formal combination results are more mixed, with a $10.15\%$ difference in formality in favor of binary and a $7.56\%$ difference in negativity in favor of GradWt. Based on these experiments, we chose to implement the GradWt approach for our remaining two- and three-style experiments, since it displays the highest overall performance on both control and generation quality.
@@ 5.2 Combinations of Two Styles
We first fine-tune RL models for all four possible combinations of sentiment and formality: positive+formal, positive+informal, negative+formal, and negative+informal. We choose these two styles since their respective discriminators have the highest F1 scores (Table 1). We use the GradWt reward formulation, as it was the most successful reward formulation based on the evaluation in Section 5.1. We consider the generations of the resulting models on our evaluation prompts, as well as the results of other approaches for multi-style controlled generations. Style accuracy results for these two-style models are in Table 4; a sample of generations produced from these models are in Appendix Table 11.
Comparison with PPLM
We extend PPLM to combine feedback from multiple discriminators (Figure  5) and find that the fine-tuned RL models have higher style control. PPLM has some success with producing more Positive+Formal and Positive+Informal generations than vanilla LLaMA2, but does not show any improvements for the Negative+Formal and Negative+Informal combinations. This suggests that PPLM may struggle to effectively combine multiple styles without incorporating some additional techniques. We leave further investigation to future work.
Comparison with Zero-Shot GPT-4
Model FineTune GradWt GPT-4
Target $\downarrow$ Sent. Form. Both Sent. Form. Both
Pos-For 0.855 0.759 0.703 0.968 0.971 0.945
Pos-Inf 0.750 0.670 0.460 0.945 0.412 0.369
Neg-Fol 0.696 0.606 0.375 0.618 0.992 0.608
Neg-Inf 0.659 0.767 0.603 0.929 0.188 0.178
Table 6: Automatic target style accuracy evaluation of 2-style models for GradWt fine-tuned LLaMA2 and GPT-4.
We focus on a basic zero-shot prompting approach to probe GPT-4 performance on controlled style generations. We chose zero-shot prompting as we are interested in scenarios with high data-scarcity, i.e. those in which the target style combination text is not readily available. We note that while few-shot prompting can out-perform zero-shot; this is not always the case – recent work suggests that performance can be equivalent or even deteriorate in a few-shot setting, e.g. Coyne et al. (2023).
We include GPT-4 results as a point of reference, but it is not a direct comparison for a few reasons. First, our other experiments use the Llama2-7B architecture, which has approximately 1,000 times fewer parameters than GPT-4. In addition, Llama2 is a base language model with no human feedback fine-tuning, unlike GPT-4. Finally, we also recognize that prompting techniques such as Chain-of-Thought prompting (Wei et al., 2022) have been shown to improve performance on many tasks and that prompts can have large impacts on GPT-4 performance; however we do not focus on prompt engineering for output optimization in this work.
@@ 5.3 Combinations of Three Styles
Styles
% Sent.
% Form.
% I/N/T
PPL
Irony, Pos, For
85.45
77.75
66.55
39.01
Neutral, Pos, For
68.65
56.20
46.65
29.19
Toxic, Neg, Inf
57.75
67.90
18.65
29.75
Neutral, Neg, Inf
55.55
73.70
52.65
39.99
Table 7: We add third style dimension such as irony and (emotional) neutrality and train a new set of three-style models. Table shows accuracy for each target style dimension (Sentiment, Formality, and third dimension Irony/Neutral/Toxic) across the 2000 generations in the evaluation set.
@@
Figure 4: Visualization of style control for two of our three-style models: Positive+Formal+Neutral (emotion) and Positive+Formal+Irony. Each style dimension shows the portion of generations containing that dimension and is min-max scaled. The two-style Positive+Formal model results are included for comparison, e.g. we see that the +Neutral model does not control formality or positivity as well as the two-style model, but the +Irony model does.
@@
The third style dimensions are from our other discriminators: toxicity, irony, and emotion. We limit our three style models to four distinct three-style combinations due to computational constraints (fully exploring the three-style combination space provided by all of our discriminators would require fine-tuning $212$ models, where each fine-tuning procedure needs to load the language model coupled with three discriminators). See Figure 4 for a visualization of these results, and Table 7 for a summary. We find that the three style models are able to increase the proportion of all three target styles relative to the base Llama2 model, and their generation quality does not decline. For some combinations (e.g. Positive-Formal-Irony), the sentiment and formality dimension control does not deteriorate compared to the two-style model. Other combinations (e.g. Negative-Informal-Toxic) show lower sentiment and formality control than the two-style models, although they still make improvements over the vanilla model. In future work, we plan to investigate how to increase policy exploration of the policy in order to encourage discovery of these more difficult style combinations.
5.4 Imbalance in Style Combinations
@@
We hypothesize that rarer style combinations are more difficult to learn. Das et al. (2023) addressed this style imbalance problem by balancing their combinatory distributions from the training data. Similarly, we approach understanding the relative frequency of the possible style combinations from two perspectives: frequency in the human texts in the datasets and frequency in the baseline generations in the datasets. In both cases, formal and negative is the least common combination, and formal and positive is the most common (see Figure 5). We observe that while our proposed method’s efficacy varies with the specific target style combination, it consistently results in more generations containing the target styles, even for difficult style combinations.
Figure 5: Percentages of evaluation generations that fulfill both target styles. Fine-tuning with GradWt-RL substantially increases prevalence of target style combinations when compared to the original base model (Llama2) or PPLM-based control.
Because the training process includes a penalty term for the KL divergence with the original language distribution in order to maintain linguistic fluency, rare style combinations that are not well-represented in the original model will likely be more difficult to produce. Figure 6 illustrates the relationship between frequency of style combinations in the human completions versus the accuracy of controlled generations for that combination.
@@
Figure 6: We compare the percentage of test generations with both target styles to the prevalence of that target style combination in the full human texts and observe a linear relationship. This suggests that more common style combinations may be easier to learn to control.
@@ 5.5 Domain-Based Performance
When considering the prompts from each of the subdatasets in the evaluation set separately, we observe that the Wikipedia portion of the evaluation set has a substantial drop in performance for Negative-Formal and Positive-Informal models with respect to target style accuracy (Appendix A Figure 7). This indicates that model performance can be strongly affected by the domain of the prompt, even in setups such as ours where the prompt only contains four words. It may also indicate that the models do not learn to generalize the target style across all domains.
@@ 6 Discussion
We propose an approach to creating generative models with multi-style control using reinforcement learning with a reward derived from a dynamic linear combination of discriminator outputs. This technique results in generations that largely conform to the target styles. However, we do observe that, especially in the 3-style set up, our approach is not successful for contradictory, rare, or difficult style combinations.
There are multiple possible approaches to the multi-style generation problem; our approach represents only one of these. In addition to PPLM, which could possibly be further optimized beyond our additions for multi-style control, other decoding-time generational control approaches include GeDI Krause et al. (2021) and DExperts Liu et al. (2021). Other fine-tuning approaches include Direct Preference Optimization (DPO) Rafailov et al. (2023). Finally, this problem can be approached via prompt engineering for LLMs such as GPT-4. Which of these approaches is best-suited to this problem with respect to accuracy, cost, and efficiency remains an open question. Future work should investigate which of these approaches results in the strongest multi-style control and their relative efficiency-accuracy trade offs.
Finally, the precise theoretical relationships between individual styles is an interesting and open question. If two styles rarely co-occur, is it because their combination is impossible, or is it simply rare? Does this distinction affect a language model’s ability to combine the two styles? Our results for the three-style models hint at the complexity of this issue. Future work should consider formalizing the feasibility of different low-frequency style combinations, and encouraging language models to explore their state space more in order to reach more rare combinations of styles.
@@ 7 Limitations
Our approach is less effective for rare style combinations like negative and formal. For control of rare style combinations, further techniques such as initial language model fine-tuning or exploration incentives are likely needed in order to achieve the desired effect. Our work also relies heavily on discriminators which may be flawed and may also be ineffective on out-of-domain inputs, limiting the interpretability of results.
We also note two side effects we observed in the generations after fine-tuning. First, there is a possibility that non-target styles may also shift as a result of the fine-tuning process. Second, the original model’s factual respresentations also appear to be altered during training.
Behavior of Uncontrolled Styles
We examine whether generations from style-controlled models differ from those of the original model with respect to uncontrolled styles. Certain textual styles are highly related, e.g. positive and joy. As expected, models targeting a positive style show a corresponding increase in the joy of their generations. However, we do observe some variations in uncontrolled styles that may not be easily explained by such associations (e.g. there is more joy in positive-formal than in positive-informal; there is more anger in negative-formal than in negative-informal. These results are summarized in Appendix Figure 8.
Factuality
We also observe that controlling for style considerably alters the model’s factual understanding. The 500-item Wikipedia subset of the evaluation set demonstrates this clearly, as the prompts (e.g. “Skydio is an American…”) lead the model to write a completion that makes factual assertions. We observe substantial fact shifting between models in these assertions, highlighting the risks of factuality deterioration from models fine-tuned for style control. See example Wikipedia generations in Appendix.

### v3

@@ 4.3 Configurations of (Custom) PPLM
Source Dataset Prompt
SemEval 2017 Its nice to stay
TweetEval @user aha but white
Rotten Tomatoes One of the most
Wikipedia Gorman Park (or Amelia
Table 2: Example prompts from each of the four source datasets used in our pipeline.
@@
Reward Formulation Negative $\uparrow$ Informal $\uparrow$ Neg & Inf $\uparrow$ PPL $\downarrow$ Bigram Dup $\downarrow$
@@
   Dynamic Weighting 65.90 76.70 60.25 31.46 0.1665
Table 3: Comparison of reward formulations for the Negative + Informal style combinations. We find that our approach (Dynamic Weighting) rewards show the best control over the style combination while maintaining generation quality.
@@ 4.4 Training and Evaluation Prompt Data
Prompts we use during training are drawn from the training sets of these datasets: SemEval emotions (Rosenthal et al., 2017), TweetEval (Barbieri et al., 2020), Rotten Tomatoes (Pang and Lee, 2005), and Wikimedia (Wikimedia, ). We choose these datasets for their variety of domains (Tweets, movie reviews, Wikipedia articles). Our evaluation prompts consist of 500 randomly selected items each datasets’ test set. Prompts consist of the first four words of each item; examples in Table 2.
4.5 Evaluation
We evaluate the generations from each model based on two criteria: first, how often the generations adhere to the target styles, and second, how well the generations maintain linguistic quality of the original model. Consider a model fine-tuned to produce positive generations: it could simply respond to every prompt with “This is great, I love it!” and force every generation to be classified as positive, but the overall language quality would suffer.
Automatic Evaluation We count the proportion of the generations that are classified by the discriminators as having the target style. We also count the proportion of the generations with both target styles to see how frequently models successfully combine styles in the same generation. To evaluate whether the generations maintain the linguistic quality of the original model, we measure the average perplexity of the generations as well as their repetitiveness (duplicate bigram frequency).
[…]
@@ 5 Results
In this section we describe our experiments for assessing the different multi-style reward formulations and the overall performance of our models for various multi-style control settings. The questions we investigate are: 1. How can we most effectively combine signals from separate style discriminators into a reward function? (§ 5.1), 2. How often do two- and three-style models express the target styles in their generation, and how fluent are these generations? (§ 5.25.3), and 3. Which style combinations are most difficult to learn? (§ 5.4).
Model Base (Llama2-7B) Dynamic Weighting (ours) PPLM
@@
Table 4: Automatic target style accuracy evaluation of 1- and 2-style models for RL and PPLM. (“Both” indicates the portion of generations containing both the target sentiment and formality.) All models (Base, Dynamic Weighting, PPLM) use Llama2-7B as the base model for fair comparison. Compare style frequencies to those in the base LLaMA2 generations to see the effect of the other two approaches.
Target Styles Inf-Neg Form-Neg
Style Lang Style Lang
Prefer Logits $8\%$ $9\%$ $8\%$ $32\%$
Prefer Dynamic Weighting (ours) 20% 69% 71% 42%
Table 5: Human evaluation results (preferences were determined by majority vote across annotators). Generations from models trained with our dynamic weighting approach were preferred with respect to both style and linguistic naturalness. (Annotators could also choose no preference.)
5.1 Multi-Reward Formulation Evaluation
We train several versions of models with Informal-Negative and Formal-Negative target styles to investigate the efficacy of the reward formulations outlined in Sec 3.2. For a summary of our study on Informal-Negative see Table 3; remaining results are in Appendix Table 7. We find that for softmax rewards, the models have minimal style control. We hypothesize that a contributing factor may be poor confidence calibration of the discriminators’ softmax scores. To address this, we implement a calibrated softmax reward using the calibration technique in Guo et al. (2017). This results in the Expected Calibration Error (ECE) for sentiment and formality decreasing from $0.133$ to $0.016$ and $0.267$ to $0.111$ respectively. These calibrated softmax rewards do offer some improvements relative to pre-calibration, but improvements are modest.
The logit reward models, on the other hand, tend to have better style control at the cost of fluency – their generations have some of the highest perplexities. During KL hyperparamater search, we also anecdotally observed that the logit models often developed highly repetitive reward hacking strategies for sub-optimal KL coeffecients.
Both the binary rewards and Dynamic Weighting rewards produce sizable improvements over other reward formulations while maintaining generation quality. Dynamic Weighting has a slight edge overall: it outperforms the binary approach on all counts for the Negative-Informal combination ($60.25\%$ of Dynamic Weighting generations contain both styles versus $56.80\%$ for binary), and its generations are less repetitive. Negative-Formal combination results are more mixed, with a $10.15\%$ difference in formality in favor of binary and a $7.56\%$ difference in negativity in favor of Dynamic Weighting. Based on these results, we chose to use the Dynamic Weighting approach for our remaining experiments, since it displays the highest overall performance on both control and generation quality.
@@ 5.1 Multi-Reward Formulation Evaluation
5.1.1 Human Evaluation
To bolster the automatic evaluation, we also conduct a human study. For cost reasons, we limit this investigation to two style combinations (negative-informal and negative-formal), and two reward formulations (Logit and Dynamic Weighting). We ask annotators whether they prefer the Logit or Dynamic Weighting completion with respect to style (randomizing the order in which the generations are displayed), and we then ask which completion they prefer with respect to the naturalness of the text. Three annotators respond to each question.
Results indicate that humans prefer the Dynamic Weighting model’s generations with respect to both style and linguistic quality (results in Table 5), echoing the conclusions of the automatic evaluations. We conclude that the automatic metrics appear to reasonably align with human perception of these aspects of the generation.
We calculated the inter-annotator agreement using Krippendorff’s alpha Krippendorff (1980). The alpha for style preference questions is 0.36 whereas for linguistic quality questions, alpha is 0.23.This indicates fair agreement among annotators, and we note that it falls in the range of alphas for subjective annotation tasks reported in Wong et al. (2021).
@@ 5.2 Combinations of Two Styles
We first fine-tune RL models for all possible combinations of sentiment and formality: positive+formal, positive+informal, negative+formal, and negative+informal. We choose these style dimensions since their respective discriminators have the highest F1 scores (Table 1). We use the Dynamic Weighting reward formulation, as it was the most successful reward formulation from Section 5.1. We consider the generations of the resulting four models on the evaluation prompts: results for these two-style models are in Table 4, and a random sample of their generations are in Appendix Table 11.
Styles % Sent. % Form. % I/N/T PPL
Irony, Pos, For 85.45 77.75 66.55 39.01
Neutral, Pos, For 68.65 56.20 46.65 29.19
Toxic, Neg, Inf 57.75 67.90 18.65 29.75
Neutral, Neg, Inf 55.55 73.70 52.65 39.99
Table 6: We add third style dimension and train a new set of three-style models. Table shows accuracy for each target style dimension (Sentiment, Formality, and third dimension Irony/Neutral (emotionally)/Toxic) across the 2000 generations in the evaluation set.
Figure 4: Percentages of evaluation generations that fulfill both target styles. Fine-tuning with Dynamic Weighting (GradWt-RL) substantially increases prevalence of target style combinations when compared to the original base model (Llama2) or PPLM-based control.
Comparison with PPLM We extend PPLM to combine feedback from multiple discriminators (Figure  4) and find that the fine-tuned RL models have higher style control. PPLM has some success with producing more Positive+Formal and Positive+Informal generations than base Llama2, but does not show any improvements for the other combinations. This suggests that PPLM may struggle to effectively combine multiple styles without incorporating some additional techniques. We leave further investigation to future work.
@@ 5.3 Combinations of Three Styles
Figure 5: Style control results for two of the 3-style models: Positive+Formal+Neutral (emotion) and Positive+Formal+Irony. Each style dimension shows the portion of generations containing that style and is min-max scaled. The two-style Positive+Formal model results are included for comparison, e.g., we see that the +Neutral model does not control formality as well as the two-style model, but the +Irony model does.
@@
The third style dimensions are from our other discriminators: toxicity, irony, and emotion. We limit our three style models to four distinct three-style combinations due to computational constraints (fully exploring the three-style combination space provided by all of our discriminators would require fine-tuning $212$ models, where each fine-tuning procedure needs to load the language model coupled with three discriminators). See Fig. 5 for a visualization of these results and Table 6 for a summary. We find that the three style models are able to increase the proportion of all three target styles relative to the base Llama2 model, and their generation quality does not decline. For some combinations (e.g., Positive-Formal-Irony), the sentiment and formality dimension control does not deteriorate compared to the two-style model. Other combinations (e.g., Negative-Informal-Toxic) show lower sentiment and formality control than the two-style models, although they still make improvements over the vanilla model.
5.4 Imbalance in Multi-Style Combinations
@@
We hypothesize that rarer style combinations are more difficult to learn. Das et al. (2023) addressed this style imbalance problem by balancing their combinatory distributions from the training data. Similarly, we approach understanding the relative frequency of the possible style combinations from two perspectives: frequency in the evaluation datasets’ baseline generations and original human texts. In both cases, formal and negative is the least common combination, while formal and positive is most common. Because the training process includes a KL divergence penalty term to maintain linguistic fluency, rare style combinations that are not well-represented in the original model will be suppressed in the controlled model’s outputs; Fig. 6 illustrates this relationship. We note that our control models are still able to improve frequencies of even the most difficult combinations (Fig. 4).
@@
Figure 6: We observe a linear relationship between the percentage of test generations containing both target styles and the prevalence of that target style combination in the full human texts. This suggests that more common style combinations may be easier to learn to control.
@@ 5.5 Domain-Based Performance
When considering the prompts from each of the subdatasets in the evaluation set separately, we observe that the Wikipedia portion of the evaluation set has a substantial drop in performance for Negative-Formal and Positive-Informal models with respect to target style accuracy (Fig. A.8). This indicates that model performance can be strongly affected by the domain of the prompt. It may also indicate that the models struggle to generalize the target style across all domains.
@@ 6 Discussion
We propose an approach to controlled multi-style generation using reinforcement learning with a reward derived from a dynamically weighted linear combination of discriminator outputs. This novel technique results in generations that largely conform to the target styles while maintaining linguistic quality.
There are multiple possible approaches to the multi-style generation problem; our approach represents only one of these. In addition to PPLM, which could possibly be further optimized beyond our additions for multi-style control, there are also the other postprocessing and retraining methods discussed in Section 2, along with other fine-tuning approaches (including Direct Preference Optimization (DPO) Rafailov et al. (2023)). In addition, this problem could be approached via prompt engineering for LLMs such as GPT-4. Which approach is best-suited to this problem with respect to accuracy, cost, and efficiency remains an open question. Future work should investigate the strongest overall approach multi-style control and the relative accuracy-efficiency trade-offs.
Finally, the precise theoretical relationships between individual styles is an interesting and open question. If two styles rarely co-occur, is it because their combination is impossible, or is it simply rare? Does this distinction affect a language model’s ability to combine the two styles? Our results for the three-style models hint at the complexity of this issue. Future work should consider formalizing the feasibility of different low-frequency style combinations, and encouraging language models to explore their state space more in order to uncover more rare combinations of styles.
@@ 7 Limitations
Our approach is less effective for rare style combinations like negative and formal. For control of rare style combinations, further techniques such as initial language model fine-tuning or exploration incentives are likely needed in order to achieve the desired effect. Our work also relies on discriminators which may be difficult to implement for data scarce attributes.
Alignment Tax
We also note two side effects we observed in the generations after fine-tuning. This is somewhat expected, as many researchers have documented performance on some benchmarking tasks decreasing as side effects of RL fine-tuning procedures (see e.g. Ouyang et al. (2022); Askell et al. (2021)). First, there is a possibility that non-target styles may also shift as a result of the fine-tuning process. Specifically, we observe some variations in uncontrolled styles after fine-tuning (e.g. there is more joy in positive-formal than in positive-informal; there is more anger in negative-formal than in negative-informal. These results are summarized in Appendix Figure 9. Second, we observe that controlling for style can alter the model’s factual claims. The 500-item Wikipedia subset of the evaluation set demonstrates this clearly, as the prompts (e.g. “Skydio is an American…”) lead the model to write a completion that makes factual assertions. See example Wikipedia generations and our annotations in Appendix A.4 Table 12. This indicates that further work is needed before such models can be relied on for factual claims.